Answers?
amortized over its useful life? would be 11. Which of the following A, $100,000 spent the B. $25,000 paid to successfully C. $150,000 paid for the exploratianof a natural resource deposit. D. $75,000 paid to purchase a regstered trademark spent on the development of a 12. Which of the following costs should be expensed when incurred? A. Replacement cost of machine component which extends the machines useful life B. Cost of a machine which will be used exclusively for research and development. C. Brokers fee for purchase of land D. Equipment delivery cost. 13. The following costs were incurred by the buyer of equipment: $10,000 875 300 500 Repair (30 days after initial use) 600 Purchase price Sales tax Insurance while in route Testing How much of those costs should be capitalized? A. $12,275 B. $11,175 C $10,875 D. $11,675 14. Poras Company purchased a machine for $80,000 on 10/1/17. The machine had an estimated useful life of 5 years and a salvage value of $5,000. Poras uses the double declining balance depreciation method. On 12/31/18, the machine was sold for $40,000. How much is Poras Companys gain or loss on the sale? A. $7,000 gain B. $3,200 loss C. $6,400 gain D. $11,200 gain 15. Edom Inc. purchased equipment for $150,000 on 7/1/17. The equipment had an estimated useful life of 4 years and a salvage value of $10,000. Edom uses the sum of method. How much depreciation should Edom record for the years digits depreciation calendar year 2018? A. $49,000 B. $42,000 C. $45,000 D. $52,500
Answer #1
11) Answer : C. $150000 paid for the exploration of a natural resource deposit.
option A,B,D are having indefinite life.
12) Answer : B. Cost of machine which will be used exclusively for research and development
All other options A,C,D, are carried forward with the asset.
13) Answer : D. $11675
Costs should be capitalized =
Pur. Price + Sales tax + Ins. In transit + testing=
10000+875+300+500=$11675
(the repair cost are revenue expense, not capital)
14) D. $11200 gain
DDB dep. Method: Rate (100/5)*2= 40%
2017 year dep. 80000*40%=32000
2018 year dep. (80000-32000)*40%=19200
Gain on sale = Cash+ Acc. Dep - Cost of purchase =
Gain on sale =40000+ (32000+19200) - 80000 = $11200
15) Answer : B. $42000
Depreciation series : 4/10,3/10,2/10&1/10
Depreciation for 2018 (2nd year) = (150000-10000)*3/10 = $42000